# Characterisation of the big Ramsey degrees of the generic partial order

### v4

@@ 4. Poset-diaries and level structures
Given a poset-diary $S$ (Definition 1.5), one can view $\overline{S}$ as a binary branching tree and each level $\overline{S}_{\ell}$ as a structure $\mathbf{S}_{\ell}=(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ where the structure $\mathbf{S}_{\ell+1}$ is constructed from the structure $\mathbf{S}_{\ell}$ as described in the following proposition.
@@
Let $S$ be a poset-diary. Then all words in $\overline{S}$ are mutually compatible, and for each $\ell\leq\sup_{w\in S}|w|$ the structures $\mathbf{S}_{\ell}=(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ and $\mathbf{S}_{\ell+1}=(\overline{S}_{\ell+1},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ are related as follows:
@@
    If $\overline{S}_{\ell}$ is splitting, then $\mathbf{S}_{\ell+1}$ is isomorphic to $\mathbf{S}_{\ell}$ with one vertex $v$ duplicated to $v^{\prime},v^{\prime\prime}$ with $v^{\prime}\mathbin{<_{\mathrm{lex}}}v^{\prime\prime}$, $v^{\prime}\not\preceq v^{\prime\prime}$, $v^{\prime\prime}\mathbin{\not\preceq}v^{\prime}$, and $v^{\prime}\trianglelefteq v^{\prime\prime}$.
@@
    It thus remains to consider the case where $u\mathbin{<_{\mathrm{lex}}}u^{\prime}$, $u,u^{\prime}\notin\{v,w\}$, are unrelated in $\overline{S}_{i}$, however their extensions are related in $\overline{S}_{i+1}$. It is not possible for $u$ to extend by $\mathrm{R}$ and $u^{\prime}$ by $\mathrm{L}$. So assume that $u$ extends by $\mathrm{X}$ and $v$ extends by $\mathrm{L}$ (the remaining case follows by symmetry). From this we conclude that $u\mathbin{<_{\mathrm{lex}}}u^{\prime}\mathbin{<_{\mathrm{lex}}}v$, $u\perp v$ and $u^{\prime}\not\perp v$. Since $u\not\perp u^{\prime}$ we again obtain a contradiction with Proposition 3.6 (P4) or (P6).
@@ 5. A poset-diary coding $\mathbf{P}$
Recall that $\mathbf{P}=(\omega,\leq_{\mathbf{P}})$ denotes a fixed enumerated generic poset. We define a function $\varphi\colon\omega\to\Sigma^{*}$ by mapping $j<\omega$ to a word $w$ of length $22j+$ 2 defined by putting $(w_{2j},w_{2j+1})=(\mathrm{L},\mathrm{R})$ and, for every $i<j$, $(w_{2i},w_{2i+1})$ to $(\mathrm{L},\mathrm{L})$ if $j\mathbin{\leq_{\mathbf{P}}}i$, $(\mathrm{R},\mathrm{R})$ if $i\mathbin{\leq_{\mathbf{P}}}j$, and $(\mathrm{X},\mathrm{X})$ otherwise. We set $T=\overline{\varphi[\omega]}$. The following result is easy to prove by induction.
Proposition 5.1 (Proposition 4.7 of [20]).
@@
$\psi^{J}_{\ell}$ is an isomorphism $(T_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)\to(\psi^{J}_{\ell}[T_{\ell}],\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$.
@@
If there are $w,w^{\prime}\in T_{\ell}$ such that $w\mathbin{<_{\mathrm{lex}}}w^{\prime}$, $w\perp w^{\prime}$ and $\psi^{J}_{\ell}(w)\mathbin{\not\perp}\psi^{J}_{\ell}(w^{\prime})$, pick among all such pairs $(w,w^{\prime})$ one minimizing |$|\{u\in T_{\ell}:w\mathbin{<_{\mathrm{lex}}}u\mathbin{\leq_{\mathrm{lex}}}w^{\prime}\}|$. Proposition 3.6 implies that the conditions in (2) are satisfied, again a contradiction with maximality of $J$.
So there are $w,w^{\prime}\in T_{\ell}$ such that $w\mathbin{<_{\mathrm{lex}}}w^{\prime}$, $w\prec w^{\prime}$ and $\psi^{J}_{\ell}(w)\mathbin{\not\prec}\psi^{J}_{\ell}(w^{\prime})$, and we can assume that $w,w^{\prime}$ maximize |$|\{u\in T_{\ell}:w\mathbin{<_{\mathrm{lex}}}u\mathbin{\leq_{\mathrm{lex}}}w^{\prime}\}|$. Proposition 3.6 implies that the conditions in (2) are satisfied, again a contradiction with maximality of $J$. Hence indeed $\psi^{J}_{\ell}$ is an isomorphism $(T_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)\to(\psi^{J}_{\ell}[T_{\ell}],\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$. ∎
@@ 6. Interesting levels and sub-diaries
We now aim to prove upper bounds for the big Ramsey degrees of $\mathbf{P}$. To do this, we need to define a notion of sub-diary which corresponds to a subtree of $\Sigma^{*}$ which preserves all important features of a given subset. Given $S\subseteq\Sigma^{*}$ we first determine which levels contain interesting changes and then define a sub-tree by removing the remaining “boring” levels from the tree. This is related to the notion of parameter-space envelopes used in [20], but more versatile, making it possible to get exact upper bounds.
@@
    the structure $\overline{\mathbf{S}}_{i}=(\overline{S}_{i},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ is not isomorphic to $\overline{\mathbf{S}}_{i+1}=(\overline{S}_{i+1},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$, or
@@
Interesting levels are the analog for subsets of $\Sigma^{*}$ of the notion of critical level for a subset of coding nodes in $\mathrm{CT}$; see for instance Definition 5.1 of [44] or Definition 5.1.3 of [3] (we note that the two definitions are slightly different).
@@
Fix $\ell\geq 0$, $S=\{u^{0}\mathbin{<_{\mathrm{lex}}}u^{1}\mathbin{<_{\mathrm{lex}}}\cdots\mathbin{<_{\mathrm{lex}}}u^{n-1}\}$, $\Sigma^{*}_{\ell}=\{v^{0}\mathbin{<_{\mathrm{lex}}}v^{1}\mathbin{<_{\mathrm{lex}}}\cdots\mathbin{<_{\mathrm{lex}}}\allowbreak v^{m-1}\}$, and a boring extension $e$ of $S$. For $u\in S$ denote by $i(u)$ the integer $i$ satisfying $u^{i}=u$. For $v\in\Sigma^{*}_{\ell}\setminus S$ we say that letter $c\in\Sigma$ is safe for $v$ if for every $0\leq j<n$ such that $u^{j}$ is compatible with $v$, it holds that ${u^{j}}^{\frown}e_{j}$ is compatible with $v^{\frown}c$ and ($(\{u^{j},v\},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ is isomorphic to ($(\{{u^{j}}^{\frown}e_{j},v^{\frown}c\},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$.
@@
Let $v^{i},v^{j}$ be compatible words. We then check that the structure ($(\{v^{i},v^{j}\},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ is isomorphic to ($(\{{v^{i}}^{\frown}e^{\prime}_{i},{v^{j}}^{\frown}e^{\prime}_{j}\},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ and that ${v^{i}}^{\frown}e^{\prime}_{i}$ and ${v^{j}}^{\frown}e^{\prime}_{j}$ are compatible. If $e^{\prime}_{i}=e^{\prime}_{j}$ (in particular, this happens whenever $v^{i}|_{\ell}=v^{j}|_{\ell}$), the result is clear.
So suppose that $e^{\prime}_{i}\neq e^{\prime}_{j}$; in particular, this implies that $v^{i}|_{\ell}\neq v^{j}|_{\ell}$. Note that in this case, the lexicographic order of $v^{i}$ and $v^{j}$ is already determined by their restrictions to level $\ell$ (that is, $v^{i}\mathbin{<_{\mathrm{lex}}}v^{j}\iff v^{i}|_{\ell}\mathbin{<_{\mathrm{lex}}}v^{j}|_{\ell}$), hence ($(\{v^{i}|_{\ell},v^{j}|_{\ell}\},\mathbin{\leq_{\mathrm{lex}}})$, ($(\{v^{i},v^{j}\},\mathbin{\leq_{\mathrm{lex}}})$, and ($(\{{v^{i}}^{\frown}e^{\prime}_{i},{v^{j}}^{\frown}e^{\prime}_{j}\},\mathbin{\leq_{\mathrm{lex}}})$ are isomorphic. Consequently, the compatibility of ${v^{i}}^{\frown}e^{\prime}_{i}$ and ${v^{j}}^{\frown}e^{\prime}_{j}$ follows from the compatibility of $v^{i}$ and $v^{j}$ and the compatibility of ${v^{i}|_{\ell}}^{\frown}e^{\prime}_{i}$ and ${v^{j}|_{\ell}}^{\frown}e^{\prime}_{j}$.
@@ 7. Upper bounds
    For all $u,v\in S$, $u\sqsubseteq v$ implies that $f(u)\sqsubseteq f(v)$.
@@
    The function $f$ is an embedding $f\colon(S,\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)\to(\Sigma^{*},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ and images of pairs of compatible words are also compatible.
@@
If $S=\Sigma^{*}_{<\ell}$ for some $\ell>0$, then the images of shape-preserving functions are always strong subtrees in the sense of Milliken’s tree theorem [41]. However, the converse is not true: For example, the function $f\colon\Sigma^{*}_{<2}\to\Sigma^{*}$ given by $f(\emptyset)=\emptyset$, $f(\mathrm{L})=\mathrm{L}\mathrm{L}$, $f(\mathrm{X})=\mathrm{X}\mathrm{R}$, and $f(\mathrm{R})=\mathrm{R}\mathrm{R}$ is not shape-preserving because all three levels 0, 1 and 2 of $f[\Sigma^{*}_{<2}]$ are interesting, while $\mathrm{Im}(f)$ is a strong subtree.
@@
Now we are finally ready to prove a big Ramsey result for $\mathbf{P}$. In the proof of the next result, it is somewhat more natural to work with colourings of copies rather than embeddings, where a copy is simply the image of an embedding. Given structures $\mathbf{A}$ and $\mathbf{B}$, we denote the set of copies of $\mathbf{A}$ in $\mathbf{B}$ by $\binom{\mathbf{B}}{\mathbf{A}}$. Note that if $\mathbf{X}\in\binom{\mathbf{B}}{\mathbf{A}}$, then there are exactly $|\mathrm{Aut}(\mathbf{A})|$-many embeddings $f\in\mathrm{Emb}(\mathbf{A},\mathbf{B})$ with $\mathrm{Im}(f)=\mathbf{X}$. From this, it follows that if one instead defines the big Ramsey degree of $\mathbf{A}$ in $\mathbf{B}$ in terms of copies rather than embeddings, the numbers will differ by exactly a factor of $|\mathrm{Aut}(\mathbf{A})|$ (see for instance Section 4 of [42]).

### v5

@@ 4. Poset-diaries and level structures
Given a poset-diary $S$ (Definition 1.5), one can view $\overline{S}$ as a binary branching tree and each level $\overline{S}_{\ell}$ as a structure $\mathbf{S}_{\ell}=(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$ where the structure $\mathbf{S}_{\ell+1}$ is constructed from the structure $\mathbf{S}_{\ell}$ as described in the following proposition.
@@
Let $S$ be a poset-diary. Then all words in $\overline{S}$ are mutually compatible, and for each $\ell\leq\sup_{w\in S}|w|$ the structures $\mathbf{S}_{\ell}=(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$ and $\mathbf{S}_{\ell+1}=(\overline{S}_{\ell+1},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$ are related as follows:
@@
    If $\overline{S}_{\ell}$ is splitting, then $\mathbf{S}_{\ell+1}$ is isomorphic to $\mathbf{S}_{\ell}$ with one vertex $v$ duplicated to $v^{\prime},v^{\prime\prime}$ with $^{\prime}\mathbin{<_{\mathrm{lex}}}v^{\prime\prime}, vv$$^{\prime}\not\preceq v^{\prime\prime}$, $v^{\prime\prime}\mathbin{\not\preceq}v^{\prime}$, and $v^{\prime}\trianglelefteq v^{\prime\prime}$.
@@
    It thus remains to consider the case where $\mathbin{<_{\mathrm{lex}}}u^{\prime}, u,u$$u^{\prime}\notin\{v,w\},$ are unrelated in $\overline{S}_{i}$, however their extensions are related in $\overline{S}_{i+1}$. It is not possible for $u$ to extend by $\mathrm{R}$ and $u^{\prime}$ by $\mathrm{L}$. So assume that $u$ extends by $\mathrm{X}$ and $v$ extends by $\mathrm{L}$ (the remaining case follows by symmetry). From this we conclude that $u\mathbin{<_{\mathrm{lex}}}u^{\prime}\mathbin{<_{\mathrm{lex}}}v$, $u\perp v$ and $u^{\prime}\not\perp v$. Since $u\not\perp u^{\prime}$ we again obtain a contradiction with Proposition 3.6 (P4) or (P6).
@@ 5. A poset-diary coding $\mathbf{P}$
Recall that $\mathbf{P}=(\omega,\leq_{\mathbf{P}})$ denotes a fixed enumerated generic poset. We define a function $\varphi\colon\omega\to\Sigma^{*}$ by mapping $j<\omega$ to a word $w$ of length $2j+2$ defined by putting $(w_{2j},w_{2j+1})=(\mathrm{L},\mathrm{R})$ and, for every $i<j$, $(w_{2i},w_{2i+1})$ to $(\mathrm{L},\mathrm{L})$ if $j\mathbin{\leq_{\mathbf{P}}}i$, $(\mathrm{R},\mathrm{R})$ if $i\mathbin{\leq_{\mathbf{P}}}j$, and $(\mathrm{X},\mathrm{X})$ otherwise. We set $T=\overline{\varphi[\omega]}$. The following result is easy to prove by induction.
Proposition 5.1 (Proposition 4.7 of [Hub25]).
@@
$\psi^{J}_{\ell}$ is an isomorphism $(T_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)\to(\psi^{J}_{\ell}[T_{\ell}],\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$.
@@
If there are $w,w^{\prime}\in T_{\ell}$ such that $\mathbin{<_{\mathrm{lex}}}w^{\prime}, ww$$\perp w^{\prime}$ and $\psi^{J}_{\ell}(w)\mathbin{\not\perp}\psi^{J}_{\ell}(w^{\prime})$, pick among all such pairs $(w,w^{\prime})$ one minimizing $|\{u\in T_{\ell}:w\mathbin{<_{\mathrm{lex}}}u\mathbin{\leq_{\mathrm{lex}}}w^{\prime}\}|$. Proposition 3.6 implies that the conditions in (2) are satisfied, again a contradiction with maximality of $J$.
So there are $w,w^{\prime}\in T_{\ell}$ such that $\mathbin{<_{\mathrm{lex}}}w^{\prime}, ww$$\prec w^{\prime}$ and $\psi^{J}_{\ell}(w)\mathbin{\not\prec}\psi^{J}_{\ell}(w^{\prime})$, and we can assume that $w,w^{\prime}$ maximize $|\{u\in T_{\ell}:w\mathbin{<_{\mathrm{lex}}}u\mathbin{\leq_{\mathrm{lex}}}w^{\prime}\}|$. Proposition 3.6 implies that the conditions in (2) are satisfied, again a contradiction with maximality of $J$. Hence indeed $\psi^{J}_{\ell}$ is an isomorphism $(T_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)\to(\psi^{J}_{\ell}[T_{\ell}],\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$. ∎
@@ 6. Interesting levels and sub-diaries
We now aim to prove upper bounds for the big Ramsey degrees of $\mathbf{P}$. To do this, we need to define a notion of sub-diary which corresponds to a subtree of $\Sigma^{*}$ which preserves all important features of a given subset. Given $S\subseteq\Sigma^{*}$ we first determine which levels contain interesting changes and then define a sub-tree by removing the remaining “boring” levels from the tree. This is related to the notion of parameter-space envelopes used in [Hub25], but more versatile, making it possible to get exact upper bounds.
@@
    the structure $\overline{\mathbf{S}}_{i}=(\overline{S}_{i},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$ is not isomorphic to $\overline{\mathbf{S}}_{i+1}=(\overline{S}_{i+1},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$, or
@@
Interesting levels are the analog for subsets of $\Sigma^{*}$ of the notion of critical level for a subset of coding nodes in $\mathrm{CT}$; see for instance Definition 5.1 of [Zuc22] or Definition 5.1.3 of [BCD+24] (we note that the two definitions are slightly different).
@@
Fix $\ell\geq 0$, $S=\{u^{0}\mathbin{<_{\mathrm{lex}}}u^{1}\mathbin{<_{\mathrm{lex}}}\cdots\mathbin{<_{\mathrm{lex}}}u^{n-1}\}$, $\Sigma^{*}_{\ell}=\{v^{0}\mathbin{<_{\mathrm{lex}}}v^{1}\mathbin{<_{\mathrm{lex}}}\cdots\mathbin{<_{\mathrm{lex}}}\allowbreak v^{m-1}\}$, and a boring extension $e$ of $S$. For $u\in S$ denote by $i(u)$ the integer $i$ satisfying $u^{i}=u$. For $v\in\Sigma^{*}_{\ell}\setminus S$ we say that letter $c\in\Sigma$ is safe for $v$ if for every $0\leq j<n$ such that $u^{j}$ is compatible with $v$, it holds that ${u^{j}}^{\frown}e_{j}$ is compatible with $v^{\frown}c$ and $(\{u^{j},v\},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$ is isomorphic to $(\{{u^{j}}^{\frown}e_{j},v^{\frown}c\},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$.
@@
Let $v^{i},v^{j}$ be compatible words. We then check that the structure $(\{v^{i},v^{j}\},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$ is isomorphic to $(\{{v^{i}}^{\frown}e^{\prime}_{i},{v^{j}}^{\frown}e^{\prime}_{j}\},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$ and that ${v^{i}}^{\frown}e^{\prime}_{i}$ and ${v^{j}}^{\frown}e^{\prime}_{j}$ are compatible. If $e^{\prime}_{i}=e^{\prime}_{j}$ (in particular, this happens whenever $v^{i}|_{\ell}=v^{j}|_{\ell}$), the result is clear.
So suppose that $e^{\prime}_{i}\neq e^{\prime}_{j}$; in particular, this implies that $v^{i}|_{\ell}\neq v^{j}|_{\ell}$. Note that in this case, the lexicographic order of $v^{i}$ and $v^{j}$ is already determined by their restrictions to level $\ell$ (that is, $v^{i}\mathbin{<_{\mathrm{lex}}}v^{j}\iff v^{i}|_{\ell}\mathbin{<_{\mathrm{lex}}}v^{j}|_{\ell}$), hence $(\{v^{i}|_{\ell},v^{j}|_{\ell}\},\mathbin{\leq_{\mathrm{lex}}})$, $(\{v^{i},v^{j}\},\mathbin{\leq_{\mathrm{lex}}})$, and $(\{{v^{i}}^{\frown}e^{\prime}_{i},{v^{j}}^{\frown}e^{\prime}_{j}\},\mathbin{\leq_{\mathrm{lex}}})$ are isomorphic. Consequently, the compatibility of ${v^{i}}^{\frown}e^{\prime}_{i}$ and ${v^{j}}^{\frown}e^{\prime}_{j}$ follows from the compatibility of $v^{i}$ and $v^{j}$ and the compatibility of ${v^{i}|_{\ell}}^{\frown}e^{\prime}_{i}$ and ${v^{j}|_{\ell}}^{\frown}e^{\prime}_{j}$.
@@ 7. Upper bounds
    For all $\in S, uvu\sqsubseteq v$ implies that $f(u)\sqsubseteq f(v)$.
@@
    The function $f$ is an embedding $f\colon(S,\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)\to(\Sigma^{*},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$ and images of pairs of compatible words are also compatible.
@@
If $S=\Sigma^{*}_{<\ell}$ for some $\ell>0$, then the images of shape-preserving functions are always strong subtrees in the sense of Milliken’s tree theorem [Tod10]. However, the converse is not true: For example, the function $f\colon\Sigma^{*}_{<2}\to\Sigma^{*}$ given by $f(\emptyset)=\emptyset$, $f(\mathrm{L})=\mathrm{L}\mathrm{L}$, $f(\mathrm{X})=\mathrm{X}\mathrm{R}$, and $f(\mathrm{R})=\mathrm{R}\mathrm{R}$ is not shape-preserving because all three levels 0, 1 and 2 of $f[\Sigma^{*}_{<2}]$ are interesting, while $\mathrm{Im}(f)$ is a strong subtree.
@@
Now we are finally ready to prove a big Ramsey result for $\mathbf{P}$. In the proof of the next result, it is somewhat more natural to work with colourings of copies rather than embeddings, where a copy is simply the image of an embedding. Given structures $\mathbf{A}$ and $\mathbf{B}$, we denote the set of copies of $\mathbf{A}$ in $\mathbf{B}$ by $\binom{\mathbf{B}}{\mathbf{A}}$. Note that if $\mathbf{X}\in\binom{\mathbf{B}}{\mathbf{A}}$, then there are exactly $|\mathrm{Aut}(\mathbf{A})|$-many embeddings $f\in\mathrm{Emb}(\mathbf{A},\mathbf{B})$ with $\mathrm{Im}(f)=\mathbf{X}$. From this, it follows that if one instead defines the big Ramsey degree of $\mathbf{A}$ in $\mathbf{B}$ in terms of copies rather than embeddings, the numbers will differ by exactly a factor of $|\mathrm{Aut}(\mathbf{A})|$ (see for instance Section 4 of [Zuc16]).